9/ Alternative compiled code sequences using instructions in classes A, B, C. Calculate the average CPI...

9/

Alternative compiled code sequences using instructions in classes A, B, C. Calculate the average CPI for sequence 1 and sequence2. Steps are required.

Class

A

B

C

CPI for class

1

3

5

IC in Sequence 1

2

4

1

IC in Sequence 2

3

1

3

10/ List the Symbols of five stages of pipelines in order.
Answer #1

Answer:

Formula for average weighted CPI = Σ i = 1 to n ( CPI * Instruction count)

Now IC in sequence 1 IC1 :

Add all sequences = 2 + 4 + 1 = 7

We know that the clock cycles = IC1 sequence * CPI for class sequences = 2*1 + 4*3 + 5 *1 = 2 + 12 + 5 = 19

Now average CPI for sequence 1 = clock cycles / IC1 = 19 / 7 = 2.7

Similarly for sequecen 2:

Add all sequence = 3 + 1 + 3 = 7

clock cycles = IC1 sequence * CPI for class sequences = 3*1 + 1*3 + 3*5 = 3 + 3 + 15 = 21

CPI for sequence 1 = clock cycles / IC2 = 21/7 = 3

Therefore average CPI for sequence 2 = 3